OpenStudy (anonymous):

A 2,100 kg car is lifted by a pulley. If the cable breaks at 4.50 m, what is the velocity of the car when it hits the ground?

OpenStudy (anonymous):

@thomaster

OpenStudy (amistre64):

looks like we might be able to use kinetic energy equation

OpenStudy (anonymous):

What is that equation.?

OpenStudy (amistre64):

\[KE=\frac12mv^2\]

OpenStudy (amistre64):

since its hit the ground then KE = PE\[PE=mgh\]

OpenStudy (amistre64):

so..\[mgh=\frac{1}{2}mv^2\] \[2gh=v^2\] finish solving for v

OpenStudy (amistre64):

there is another route we could take, but this ones fair enough

OpenStudy (anonymous):

Im not sure if i did it right but I got 88.2

OpenStudy (amistre64):

lets verify sqrt(2(9.8)(4.5)) ... you didnt sqrt it

OpenStudy (anonymous):

So it would be 9.39 right?

OpenStudy (anonymous):

and the units would be m/s

OpenStudy (amistre64):

another method is simply to use the formula:\[s=\frac12at^2+v_ot+h_o\]and solve for t to determine the solution to:\[v=at+v_o\] ----------------------------- \[0=-\frac12(9.8)t^2+4.5\] \[9=9.8t^2\] \[\sqrt{\frac{9}{9.8}}=t\] \[v=9.8\sqrt{\frac{9}{9.8}}\] \[v=\sqrt{9(9.8)}\] \[v=\sqrt{2(4.5)(9.8)}\] yay its the same

OpenStudy (amistre64):

yes, 9.39 m/s

OpenStudy (anonymous):

Okay :D thank you and are you good at refraction?

OpenStudy (amistre64):

i dont have any experience with refraction so prolly not

OpenStudy (anonymous):

Hmm okay well thank you :D

OpenStudy (amistre64):

youre welcome, and good luck :)